# MCP-SafetyBench: A Benchmark for Safety Evaluation of Large Language Models with Real-World MCP Servers

### v1

@@ Abstract
Large language models (LLMs) are evolving into agentic systems that reason, plan, and operate external tools. The Model Context Protocol (MCP) is a key enabler of this transition, offering a standardized interface for connecting LLMs with heterogeneous tools and services. Yet MCP’s openness and multi-server workflows introduce new safety risks that existing benchmarks fail to capture, as they focus on isolated attacks or lack real-world coverage. We present MCP-SafetyBench, a comprehensive benchmark built on real MCP servers that supports realistic multi-turn evaluation across five domains—browser automation, financial analysis, location navigation, repository management, and web search. It incorporates a unified taxonomy of 20 MCP attack types spanning server, host, and user sides, and includes tasks requiring multi-step reasoning and cross-server coordination under uncertainty. Using MCP-SafetyBench, we systematically evaluate leading open- and closed-source LLMs, revealing large disparities in safety performance and escalating vulnerabilities as task horizons and server interactions grow. Our results highlight the urgent need for stronger defenses and establish MCP-SafetyBench as a foundation for diagnosing and mitigating safety risks in real-world MCP deployments. Our benchmark is available at https://github.com/xjzzzzzzzz/MCPSafety.
@@ 1 Introduction
Large language models (LLMs) are rapidly evolving from passive text generators (brown2020languagemodelsfewshotlearners; ouyang2022traininglanguagemodelsfollow) into agentic systems capable of reasoning, planning, and operating external tools (deepseekai2025deepseekr1incentivizingreasoningcapability; react; GLM4_5; KimiK2; GPT5; anthropic2024mcp). A key driver of this shift is the Model Context Protocol (MCP) (anthropic2024mcp), which standardizes how LLMs connect to tools, data sources, and services. By abstracting and unifying API calls, MCP enables agents to dynamically discover and invoke tools across heterogeneous servers and environments. This design greatly reduces integration complexity and has accelerated the widespread adoption of MCP in both academia and industry (OpenAIMCP; ClineMCP; CursorMCP; GoogleMCP).
@@
However, the openness and extensibility of MCP introduce new safety risks (hou2025modelcontextprotocolmcp). For example, attackers can embed malicious instructions in tool metadata or descriptions, misleading models during tool invocation (kellner_fischer_2025). Attackers can also poison context during cross-server propagation (e.g., context poisoning), leading to persistent chain contamination (croce2025trivial). Moreover, malicious servers with high privileges can trigger unauthorized actions or exfiltrate sensitive data (jing2025mcipprotectingmcpsafety). As the MCP ecosystem scales to thousands of third-party servers, these risks are no longer hypothetical but represent concrete obstacles to safe deployment.
@@
Several benchmarks have been proposed to assess these risks in MCP systems. While existing MCP safety benchmarks such as SHADE-Arena (kutasov2025shadearenaevaluatingsabotagemonitoring), SafeMCP (fang2025identifymitigatethirdpartysafety), MCPTox (wang2025mcptoxbenchmarktoolpoisoning), MCIP-Bench (jing2025mcipprotectingmcpsafety), MCP-AttackBench (xing2025mcpguarddefenseframeworkmodel), and MCPSecBench (yang2025mcpsecbenchsystematicsecuritybenchmark) have provided valuable foundations for studying MCP attacks, most of them either focus narrowly on specific attack types or lack integration with realistic MCP servers. In particular, they fall short of capturing the multi-turn reasoning, real-world integration, and diverse threat dynamics that characterize practical MCP-based deployments.
@@
In this paper, we present MCP-SafetyBench, a comprehensive benchmark designed to systematically evaluate the robustness of LLM Agents against MCP attacks. Built on the MCP-Universe benchmark (luo2025mcpuniversebenchmarkinglargelanguage), MCP-SafetyBench provides tasks that reflect realistic scenarios and multi-turn reasoning workflows, filling critical gaps in existing evaluations. Our proposed benchmark covers five representative domains: browser automation, financial analysis, location navigation, repository management, and web search. It further encompasses 20 distinct attack types across the MCP server, host, and user sides. Unlike prior benchmarks limited to one-shot tool use, MCP-SafetyBench captures the inherently multi-turn nature of real-world scenarios, where attacks can emerge at any step of the interaction. Figure 2 illustrates a attack Tool Poisoning – Parameter Poisoning case. The user requests JNJ holdings, but the tool manifest silently rewrites the ticker to TSLA, causing the agent to plan correctly yet execute on the wrong target. The task is marked Fail by the task evaluator and Success by the attack evaluator, demonstrating how MCP-SafetyBench exposes hidden vulnerabilities in realistic multi-turn MCP-based workflows. The tasks in our MCP-SafetyBench are real-world tasks, requiring models to perform multi-step reasoning and coordinate across multiple servers under uncertain conditions. MCP-SafetyBench further includes 20 distinct attack types spanning the MCP server, MCP host, and user levels.
@@
We conduct a systematic evaluation of both open-source and proprietary LLMs on MCP-SafetyBench. Figure 1 shows a clear negative trend between Defense Success Rate and Task Success Rate, indicating that no model achieves both strong task performance and robust defense. In our experiment section, we further reveal substantial disparities in safety resilience, with vulnerabilities compounding as the scope of tasks expands and server interactions become more complex. These results demonstrate that MCP agents face serious and escalating safety risks, underscoring the urgent need for stronger defenses. MCP-SafetyBench thus provides a solid foundation for diagnosing safety challenges in the rapidly expanding MCP ecosystem.
@@
In summary, our contributions are as follows: 1) We develop a unified taxonomy of 20 MCP attack types that consolidates prior work and clarifies key attack categories; 2) We build MCP-SafetyBench, a benchmark based on this taxonomy and real-world MCP servers, supporting realistic multi-step safety evaluation across five domains; 3) We systematically evaluate leading open-source and proprietary LLMs, revealing large differences in safety performance and escalating vulnerabilities in multi-turn, multi-server settings.
@@
SafeMCP (fang2025identifymitigatethirdpartysafety) × × × 2 10
MCPTox (wang2025mcptoxbenchmarktoolpoisoning) × × × 11 45
MCIP-bench (jing2025mcipprotectingmcpsafety) × × × 10 -
MCP-AttackBench (xing2025mcpguarddefenseframeworkmodel) × × × 10 -
MCPSecBench (yang2025mcpsecbenchsystematicsecuritybenchmark) × × 17 -
@@ 2 Related Work
Model Context Protocol. The Model Context Protocol (MCP), introduced by Anthropic in late 2024, standardizes interaction between AI agents and external tools (anthropic2024mcp). Built on JSON-RPC 2.0 over STDIO and SSE, MCP addresses the long-standing “data silo” problem by allowing agents to dynamically discover, select, and orchestrate tools according to task context (Edwin2025). It adopts a three-layer architecture: Host (LLM agent), Client (message bridge managing user interaction), and Servers (tool and resource providers) (anthropic2024mcp). In practice, the Host connects to Servers via the Client, registers tool metadata, invokes tools, and synthesizes results into final outputs.
@@
MCP Server
Tool Poisoning-Parameter Poisoning × × Modify tool parameters in tool descriptions
@@
Function Overlapping × × × Malicious tools resemble trusted ones
@@
MCP Host
Intent Injection × × × Alter user intent
@@
User
Malicious Code Execution × × × × User input causes tools to execute harmful commands
@@
Attack Vectors in MCP. While MCP extends agent capability, it also creates new security exposures. Invariant Labs (kellner_fischer_2025) introduced Tool Poisoning Attacks (TPA), where malicious metadata or instructions inside tool descriptions manipulate agent behavior. They further proposed Shadowing Attacks, in which malicious servers override trusted tools, and MCP Rug Pulls, where benign tools are later updated with harmful logic. Follow-up work broadened the surface:  wang2025mpmapreferencemanipulationattack demonstrated Preference Manipulation Attacks, biasing tool choice through persuasive descriptions.  radosevich2025mcpsafetyauditllms proposed four attack types: malicious code execution, remote access control, credential theft, and retrieval-agent deception.  hou2025modelcontextprotocolmcp organized these risks into a lifecycle taxonomy spanning configuration, interaction, and termination phases, covering threats like name collisions, function overlapping, and sandbox escape. Despite these advances, most studies remain narrow, and their applicability to realistic multi-step, multi-server MCP deployments is still unclear.
@@
MCP Safety Benchmarks. Recent benchmarks explore MCP security from different angles. SHADE-Arena (kutasov2025shadearenaevaluatingsabotagemonitoring) studies sabotage behaviors in virtual environments. SafeMCP (fang2025identifymitigatethirdpartysafety) evaluates third-party service risks with both passive and active defenses. MCPTox (wang2025mcptoxbenchmarktoolpoisoning) focuses on tool poisoning vulnerabilities. MCIP-bench (jing2025mcipprotectingmcpsafety) builds a taxonomy-driven dataset from function-calling corpora, and MCP-AttackBench (xing2025mcpguarddefenseframeworkmodel) scales adversarial testing with 70k+ attack samples. Beyond MCP-specific settings, MCPSecBench (yang2025mcpsecbenchsystematicsecuritybenchmark) provides a modular testbed covering seventeen representative attacks across user, host, transport, and server layers.
Table 1 compares these benchmarks in attack coverage, domains, and evaluation settings. In contrast, our MCP-SafetyBench centers on real-world MCP servers, supports multi-step and multi-server interactions, and covers twenty attack types across five domains for a more realistic and comprehensive safety evaluation.
@@
Positioning within Broader Safety Frameworks. Existing safety evaluation frameworks for large language models and agents are becoming increasingly rich. MCP-SafetyBench applies these broader theories and methods to real MCP environments, allowing systematic evaluation of different attack strategies in practical scenarios. It instantiates threats in the deployment-stage from OTM (verma2025otm) and maps its taxonomy to MCP-specific threats: user-side attacks correspond to OTM’s Application Input Layer, server-side attacks correspond to the Context Data Layer, and host-side vulnerabilities correspond to internal logic attacks. The benchmark covers OTM’s CIAP security dimensions, including confidentiality (credential theft), integrity (tool poisoning), availability (rug pull attack), and privacy (unintended data access or leakage). In addition, MCP-SafetyBench complements evaluations of agent behavior and tool calling (feffer2024redteaminggenerativeaisilver; nakash2024breakingreactagentsfootinthedoor), focusing on actual risks during task execution rather than static prompt-based tests.
@@ 3.1 Overview
To address the gap in realistic safety evaluation for LLM agents, we introduce MCP-SafetyBench, a comprehensive benchmark designed to evaluate the robustness of LLM agents interacting with real MCP servers in multi-step, tool-using tasks. Unlike prior one-shot or simulated evaluations (fang2025identifymitigatethirdpartysafety; wang2025mcptoxbenchmarktoolpoisoning; jing2025mcipprotectingmcpsafety; xing2025mcpguarddefenseframeworkmodel; yang2025mcpsecbenchsystematicsecuritybenchmark), it targets security risks in ReAct-style agents (react) and it is built upon three core principles: realism, ensuring tasks mirror real-world applications; coverage, systematically targeting vulnerabilities across the entire MCP stack; and reproducibility, enabling deterministic, execution-based evaluation. Our proposed benchmark enables systematic assessment along two key dimensions: task success, which measures whether the user’s goal is achieved, and attack success, which determines if the attacker’s objective is realized, either through disruption or stealth.
@@ 3.2 MCP Attack Taxonomy
MCP Server-Side Attacks Scope. Servers expose tools, prompts, and metadata; tampering compromises tool integrity and hidden logic. Representative Types. Tool Poisoning (parameter, command, filesystem, redirection, network, dependency); Function Overlapping; Preference Manipulation; Tool Shadowing; Function Return Injection; Rug Pull
@@ 3.3 Benchmark Design and Construction
MCP-SafetyBench is constructed through a three-stage process that transforms standard tasks from the MCP-Universe benchmark (luo2025mcpuniversebenchmarkinglargelanguage) into robust security test cases. This process yields attack-instrumented tasks across five domains, including browser automation, financial analysis, location navigation, repository management, and web search. Each task is paired with exactly one MCP-layer attack drawn from our taxonomy, enabling controlled evaluation of both correctness and security outcomes. As shown in Figure 3, the construction process involves three steps:
@@ 4.1 Experiment settings
Agent Framework. In our experiments, we adopt the ReAct framework (react), which has proven to be one of the most widely used and time-tested paradigms for building agentic LLM systems. ReAct allows models to interleave reasoning and acting, making it particularly suitable for the multi-step, multi-server MCP environment.
Models. Our evaluation includes a representative set of state-of-the-art proprietary and open-source Large Language Models. For proprietary models, we include OpenAI’s GPT-5  (GPT5), GPT-4.1 (GPT4-1), GPT-4o (GPT4o), o4-mini  (o3-o4mini), Anthropic’s Claude-4.0-Sonnet (claude4) and Claude-3.7-Sonnet (claude3.7), Google’s Gemini-2.5-Pro (Gemini2-5) and Gemini-2.5-Flash (Gemini2-5), and xAI’s Grok-4 (grok4). For open-source models, we consider Zhipu’s GLM-4.5 (GLM4_5), Moonshot’s Kimi-K2 (KimiK2), Qwen’s Qwen3-235B (Qwen3), and DeepSeek’s DeepSeek-V3.1 (DeepSeek-V3).

### v2

@@ Abstract
Large language models (LLMs) are evolving into agentic systems that reason, plan, and operate external tools. The Model Context Protocol (MCP) is a key enabler of this transition, offering a standardized interface for connecting LLMs with heterogeneous tools and services. Yet MCP’s openness and multi-server workflows introduce new safety risks that existing benchmarks fail to capture, as they focus on isolated attacks or lack real-world coverage. We present MCP-SafetyBench, a comprehensive benchmark built on real MCP servers that supports realistic multi-turn evaluation across five domains—browser automation, financial analysis, location navigation, repository management, and web search. It incorporates a unified taxonomy of 20 MCP attack types spanning server, host, and user sides, and includes tasks requiring multi-step reasoning and cross-server coordination under uncertainty. Using MCP-SafetyBench, we systematically evaluate leading open- and closed-source LLMs, revealing that all models remain vulnerable to MCP attacks, with a notable safety-utility trade-off. Our results highlight the urgent need for stronger defenses and establish MCP-SafetyBench as a foundation for diagnosing and mitigating safety risks in real-world MCP deployments. Our benchmark is available at https://github.com/xjzzzzzzzz/MCPSafety.
@@ 1 Introduction
Large language models (LLMs) are rapidly evolving from passive text generators (Brown et al., 2020; Ouyang et al., 2022) into agentic systems capable of reasoning, planning, and operating external tools (DeepSeek-AI et al., 2025a; Yao et al., 2023; Zai, 2025; Moonshot, 2025; OpenAI, 2025b; Anthropic, 2024). A key driver of this shift is the Model Context Protocol (MCP) (Anthropic, 2024), which standardizes how LLMs connect to tools, data sources, and services. By abstracting and unifying API calls, MCP enables agents to dynamically discover and invoke tools across heterogeneous servers and environments. This design greatly reduces integration complexity and has accelerated the widespread adoption of MCP in both academia and industry (OpenAI, ; Cline, ; Cursor, ; Google, ).
@@
However, the openness and extensibility of MCP introduce new safety risks (Hou et al., 2025). For example, attackers can embed malicious instructions in tool metadata or descriptions, misleading models during tool invocation (Beurer-Kellner and Fischer, 2025). Attackers can also poison context during cross-server propagation (e.g., context poisoning), leading to persistent chain contamination (Croce and South, 2025). Moreover, malicious servers with high privileges can trigger unauthorized actions or exfiltrate sensitive data (Jing et al., 2025). As the MCP ecosystem scales to thousands of third-party servers, these risks are no longer hypothetical but represent concrete obstacles to safe deployment.
@@
Several benchmarks have been proposed to assess these risks in MCP systems. While existing MCP safety benchmarks such as SHADE-Arena (Kutasov et al., 2025), SafeMCP (Fang et al., 2025), MCPTox (Wang et al., 2025a), MCIP-Bench (Jing et al., 2025), MCP-AttackBench (Xing et al., 2025), and MCPSecBench (Yang et al., 2025b) have provided valuable foundations for studying MCP attacks, most of them either focus narrowly on specific attack types or lack integration with realistic MCP servers. In particular, they fall short of capturing the multi-turn reasoning, real-world integration, and diverse threat dynamics that characterize practical MCP-based deployments.
@@
In this paper, we present MCP-SafetyBench, a comprehensive benchmark designed to systematically evaluate the robustness of LLM Agents against MCP attacks. Built on the MCP-Universe benchmark (Luo et al., 2025), MCP-SafetyBench provides tasks that reflect realistic scenarios and multi-turn reasoning workflows, filling critical gaps in existing evaluations. Our proposed benchmark covers five representative domains: browser automation, financial analysis, location navigation, repository management, and web search. It further encompasses 20 distinct attack types across the MCP server, host, and user sides. Unlike prior benchmarks limited to one-shot tool use, MCP-SafetyBench captures the inherently multi-turn nature of real-world scenarios, where attacks can emerge at any step of the interaction. Figure 2 illustrates a attack Tool Poisoning – Parameter Poisoning case. The user requests JNJ holdings, but the tool manifest silently rewrites the ticker to TSLA, causing the agent to plan correctly yet execute on the wrong target. The task is marked Fail by the task evaluator and Success by the attack evaluator, demonstrating how MCP-SafetyBench exposes hidden vulnerabilities in realistic multi-turn MCP-based workflows. The tasks in our MCP-SafetyBench are real-world tasks, requiring models to perform multi-step reasoning and coordinate across multiple servers under uncertain conditions. MCP-SafetyBench further includes 20 distinct attack types spanning the MCP server, MCP host, and user levels.
@@
We conduct a systematic evaluation of both open-source and proprietary LLMs on MCP-SafetyBench. Figure 1 shows a clear negative trend between Defense Success Rate and Task Success Rate, indicating that no model achieves both strong task performance and robust defense. In our experiment section, we further reveal substantial disparities in safety resilience and variable attack effectiveness—with host-side attacks achieving the highest attack success rate. These results demonstrate that MCP agents face serious and escalating safety risks, underscoring the urgent need for stronger defenses. MCP-SafetyBench thus provides a solid foundation for diagnosing safety challenges in the rapidly expanding MCP ecosystem.
@@
In summary, our contributions are as follows: 1) We develop a unified taxonomy of 20 MCP attack types that consolidates prior work and clarifies key attack categories; 2) We build MCP-SafetyBench, a benchmark based on this taxonomy and real-world MCP servers, supporting realistic multi-step safety evaluation across five domains; 3) We systematically evaluate leading open-source and proprietary LLMs, revealing large differences in safety performance across models and variable effectiveness among different attack types.
@@
SafeMCP (Fang et al., 2025) × × × 2 10
MCPTox (Wang et al., 2025a) × × × 11 45
MCIP-bench (Jing et al., 2025) × × × 10 -
MCP-AttackBench (Xing et al., 2025) × × × 10 -
MCPSecBench (Yang et al., 2025b) × × 17 -
@@ 2 Related Work
Model Context Protocol. The Model Context Protocol (MCP), introduced by Anthropic in late 2024, standardizes interaction between AI agents and external tools (Anthropic, 2024). Built on JSON-RPC 2.0 over STDIO and SSE, MCP addresses the long-standing “data silo” problem by allowing agents to dynamically discover, select, and orchestrate tools according to task context (Edwin, 2025). It adopts a three-layer architecture: Host (LLM agent), Client (message bridge managing user interaction), and Servers (tool and resource providers) (Anthropic, 2024). In practice, the Host connects to Servers via the Client, registers tool metadata, invokes tools, and synthesizes results into final outputs.
@@
MCP Server Tool Poisoning-Parameter Poisoning × × Modify tool parameters in tool descriptions
@@
Function Overlapping × × × Malicious tools use similar names to trusted ones
@@
MCP Host Intent Injection × × × Alter user intent
@@
User Malicious Code Execution × × × × User input causes tools to execute harmful commands
@@
Attack Vectors in MCP. While MCP extends agent capability, it also creates new security exposures (Guo et al., 2025). Invariant Labs (Beurer-Kellner and Fischer, 2025) introduced Tool Poisoning Attacks, where malicious metadata or instructions inside tool descriptions manipulate agent behavior. They further proposed Shadowing Attacks, in which malicious servers override trusted tools, and MCP Rug Pulls, where benign tools are later updated with harmful logic. Follow-up work broadened the surface:  Wang et al. (2025b) demonstrated Preference Manipulation Attacks, biasing tool choice through persuasive descriptions.  Radosevich and Halloran (2025) proposed four attack types: Malicious Code Execution, Remote Access Control, Credential Theft, and Retrieval-Agent Deception.  Hou et al. (2025) organized these risks into a lifecycle taxonomy, covering threats like Tool Name Conflict and Sandbox Escape.  Jing et al. (2025) further categorized these risks by source, scope, and type, encompassing client-side attacks: Intent Injection, Data Injection, Identity Injection, and Replay Injection. Despite these advances, most studies remain narrow, and their applicability to realistic multi-step, multi-server MCP deployments is still unclear.
@@
MCP Safety Benchmarks. Recent benchmarks explore MCP security from different angles. SHADE-Arena (Kutasov et al., 2025) studies sabotage behaviors in virtual environments. SafeMCP (Fang et al., 2025) evaluates third-party service risks with both passive and active defenses. MCPTox (Wang et al., 2025a) focuses on tool poisoning vulnerabilities. MCIP-bench (Jing et al., 2025) builds a taxonomy-driven dataset from function-calling corpora, and MCP-AttackBench (Xing et al., 2025) scales adversarial testing with 70k+ attack samples. MCPSecBench (Yang et al., 2025b) provides a benchmark covering seventeen representative attacks across user, host, transport, and server layers. In contrast, our MCP-SafetyBench centers on real-world MCP servers, supports multi-step and multi-server interactions, and covers twenty attack types across five domains for a more realistic and comprehensive safety evaluation. Table 1 compares these benchmarks in attack coverage, domains, and evaluation settings.
@@
Positioning within Broader Safety Frameworks. Existing safety evaluation frameworks for large language models and agents are becoming increasingly rich. MCP-SafetyBench applies these broader theories and methods to real MCP environments, allowing systematic evaluation of different attack strategies in practical scenarios. It instantiates threats in the deployment-stage from OTM (Verma et al., 2024) and maps its taxonomy to MCP-specific threats: user-side attacks correspond to OTM’s Application Input Layer, server-side attacks correspond to the Context Data Layer, and host-side vulnerabilities correspond to internal logic attacks. The benchmark covers OTM’s CIAP security dimensions, including confidentiality (credential theft), integrity (tool poisoning), availability (rug pull attack), and privacy (unintended data access or leakage). In addition, MCP-SafetyBench complements evaluations of agent behavior and tool calling (Feffer et al., 2024; Nakash et al., 2025), focusing on actual risks during task execution rather than static prompt-based tests.
@@ 3.1 Overview
To address the gap in realistic safety evaluation for LLM agents, we introduce MCP-SafetyBench, a comprehensive benchmark designed to evaluate the robustness of LLM agents interacting with real MCP servers in multi-step, tool-using tasks. Unlike prior one-shot or simulated evaluations (Fang et al., 2025; Wang et al., 2025a; Jing et al., 2025; Xing et al., 2025; Yang et al., 2025b), it targets security risks in ReAct-style agents (Yao et al., 2023) and it is built upon three core principles: realism, ensuring tasks mirror real-world applications; coverage, systematically targeting vulnerabilities across the entire MCP stack; and reproducibility, enabling deterministic, execution-based evaluation. Our proposed benchmark enables systematic assessment along two key dimensions: task success, which measures whether the user’s goal is achieved, and attack success, which determines if the attacker’s objective is realized, either through disruption or stealth.
@@ 3.2 MCP Attack Taxonomy
MCP Server-Side Attacks Scope. Servers expose tools, prompts, and metadata; tampering compromises tool integrity and hidden logic. Representative Types. Tool Poisoning (parameter, command, filesystem, redirection, network, dependency); Function Overlapping; Preference Manipulation; Tool Shadowing; Function Return Injection; Rug Pull Attack
@@ 3.3 Benchmark Design and Construction
MCP-SafetyBench is constructed through a three-stage process that transforms standard tasks from the MCP-Universe benchmark (Luo et al., 2025) into robust security test cases. This process yields attack-instrumented tasks across five domains, including browser automation, financial analysis, location navigation, repository management, and web search. Each task is paired with exactly one MCP-layer attack drawn from our taxonomy, enabling controlled evaluation of both correctness and security outcomes. As shown in Figure 3, the construction process involves three steps:
@@ 4.1 Experiment settings
Agent Framework. In our experiments, we adopt the ReAct framework (Yao et al., 2023), which has proven to be one of the most widely used and time-tested paradigms for building agentic LLM systems. ReAct allows models to interleave reasoning and acting, making it particularly suitable for the multi-step, multi-server MCP environment.
Models. Our evaluation includes a representative set of state-of-the-art proprietary and open-source Large Language Models. For proprietary models, we include OpenAI’s GPT-5  (OpenAI, 2025b), GPT-4.1 (OpenAI, 2025a), GPT-4o (OpenAI et al., 2024), o4-mini  (OpenAI, 2025c), Anthropic’s Claude-4.0-Sonnet (Anthropic, 2025b) and Claude-3.7-Sonnet (Anthropic, 2025a), Google’s Gemini-2.5-Pro (Comanici et al., 2025) and Gemini-2.5-Flash (Comanici et al., 2025), and xAI’s Grok-4 (xAI, 2025). For open-source models, we consider Zhipu’s GLM-4.5 (Zai, 2025), Moonshot’s Kimi-K2 (Moonshot, 2025), Qwen’s Qwen3-235B (Yang et al., 2025a), and DeepSeek’s DeepSeek-V3.1 (DeepSeek-AI et al., 2025b).